OpenStudy (anonymous):

-2 (x - 2) = 6 what is the process for the correct answer

OpenStudy (anonymous):

-2(x-2) = 6 -2x+4 = 6 -2x = 2 x = -1

OpenStudy (anonymous):

-(x-2)= 6/2 -x + 2 = 3 -x = 3 - 2 x = -1

OpenStudy (anonymous):

m/4 = 13 help!

OpenStudy (anonymous):

m = 4* 13